Using Appendix C-1 or Appendix C-2 find the p-value for each test statistic. (Round your answers to 4 decimal places.) Test Statistic z = +1.34 p-value (a) Right-tailed test (b) Left-tailed test (c) Two-tailed test z-_2.07 z =-1.69
Answer #1

a)

For right tailed test p value is calculated as

P( Z > 1.34) = 1 - P( Z < 1.34)

= 1 - 0.9099

= 0.0901

p value = 0.0901

b)

For left tailed test p value is calculated as

P( Z < -2.07) = 1 - p( Z < 2.07)

= 1 - 0.9808

= 0.0192

p value = 0.0192

c)

For ltwo tailed test p value is calculated as

p value = 2 * P( Z < 1.69)

= 2 * ( 1 - P( Z < 1.69) )

= 2 * ( 1 - 0.0455)

= 0.0910